OpenStudy (karolynxx):

Which best describes the relationship between the successive terms in the sequence shown? 9, –1, –11, –21, … a)The common difference is –10. b)The common difference is 10. c)The common ratio is –9. d)The common ratio is 9. I believe i understand it but have to ask it would be the common difference is -10 bevause you would take -1-9 and get -10?

OpenStudy (karolynxx):

@freckles

OpenStudy (freckles):

if the sequence has a common difference then you can find that common difference by doing term-previous term just as if the sequence had a common ratio then you can find that common ratio by doing term divided by previous term =term/previous so since you are saying it has a common difference that means you are saying all the following differences are the same: term-previous -1-9 -11-(-1) -21-(-11)

OpenStudy (freckles):

which is true -10 is the difference for all of those

OpenStudy (freckles):

you are right! :)

OpenStudy (karolynxx):

Finally i believe im understanding .

OpenStudy (freckles):

yah!

OpenStudy (karolynxx):

A sequence is defined recursively using the formula f(n + 1) = f(n) – 5 . Which sequence could be generated using this formula? a)1, –5, 25, –125, ... b)2, 10, 50, 250, ... c)3, –2, –7, –12, ... d)4, 9, 14, 19, ... so in this equation i am subtracting -5 everytime?

OpenStudy (freckles):

yes to find the next term

OpenStudy (karolynxx):

so it would be C right ?

OpenStudy (freckles):

3 3-5=-2 -2-5=-7 -7-5=-12 that works

OpenStudy (karolynxx):

Thank you!

OpenStudy (freckles):

what if it said f(n+1)=f(n)+5

OpenStudy (karolynxx):

you would add!

OpenStudy (freckles):

right and which choice would it have been

OpenStudy (karolynxx):

so if a equation that would work would be like 2,7,12,17..

OpenStudy (freckles):

well I was talking about your choices you just posted but yes that one follows the same form for f(n+1)=f(n)+5 a)1, –5, 25, –125, ... b)2, 10, 50, 250, ... c)3, –2, –7, –12, ... d)4, 9, 14, 19, ... do you see that a would be f(n+1)=-5f(n) and that b would be f(n+1)=5(n)

OpenStudy (karolynxx):

oh the choices up there it would have been D!

OpenStudy (freckles):

yep yep :)

OpenStudy (karolynxx):

B is f(n) because thats multiplying correct ?

OpenStudy (freckles):

oops I made a type-o *and that b would be f(n+1)=5f(n)* that f part is very important

OpenStudy (karolynxx):

i see the difference Thanks !

OpenStudy (freckles):

cool stuff I kind of like that they had all of those has choices

OpenStudy (freckles):

as choices*

OpenStudy (karolynxx):

me too , i liked how you broke it down that each anwser choice had its own anwser to a equation .

OpenStudy (freckles):

it might not always be like that but this one was definitely like that

OpenStudy (karolynxx):

i understand that , but thats whats useful about understanig when to multiply add or subtract.

OpenStudy (freckles):

agreed

OpenStudy (freckles):

a)1, –5, 25, –125, ... b)2, 10, 50, 250, ... c)3, –2, –7, –12, ... d)4, 9, 14, 19, ... summary of that last question just so the thoughts can be gathered up into one post a) has form f(n+1)=-5f(n) b) has form f(n+1)=5f(n) c) has form f(n+1)=f(n)-5 d) has form f(n+1)=f(n)+5

OpenStudy (karolynxx):

Makes alot more sense like that!

OpenStudy (freckles):

when you see something like f(n+1)=rf(n) think multiplication by r to get next term when you see something like f(n+1)=f(n)+d think addition by d to get next term

OpenStudy (freckles):

anyways I think I'm going to go to bed now good luck studying geometric and arithmetic recursive style